Grade 11Physical Chemistry

A sample of clay contains 40% silica and 15% water. The sample is partially dried by which it loses 5g water. If the percentage of water in the partially dried clay is 8, calculate the percentage of silica in the partially dried clay.

Mayur Reddy
10 Years agoGrade 11

1 Answer

Gaurav Rathod
8 Years ago
Clay original sample contains 40% silica , 15% water so total non water component is 100 - 15 = 85 gram .Now sample is partially dried and loses 5 gram water .Again dried clay contains 8 % water means 8 g water . Therefore total non water component is 100-8 = 92 gram.Now percent of silica in partially dried sample will be 40/85×92= 43.29%